A container holds 50.0mL of nitrogen at 25C and a pressure of 736mm Hg. What will be it’s volume if the temperature increases by 35C ?
1
Expert's answer
2020-04-10T13:37:50-0400


A container holds 50.0mL of nitrogen at 25C and a pressure of 736mm Hg. What will be it’s volume if the temperature increases by 35C ?


Solution:

According to Mendeleev-Clapeyron's equation:

pV = nRT,

where p = 736 mm Hg = 1.0006 atm; T = 273 + 25 = 298 K; V = 50.0 mL = 0.050 L; R = 0.082 - universal gas constant.

So number of moles of N2 molecule equal:

n = pV/(RT) = 1.0006*0.050/(0.082*298) = 0.002 (moles)

Using Mendeleev-Clapeyron's equation, we can determine the volume if the temperature increases by 35C:

V2 = nRT2/p = 0.002*0.082*(273+35)/1.0006 = 0.0505 (L) = 50.5 (mL)


Answer: V2 = 50.5 mL.